# Low-latency Federated LLM Fine-tuning Over Wireless Networks

### v1

@@ I Introduction
Although integrating FL and LLMs offers a promising solution to collaboratively fine-tune a shared LLM [18], two major challenges remain. First, the massive scale of LLMs imposes heavy computation and communication demands, which hinders deployment in realistic edge device environments [17]. Second, clients with limited and heterogeneous computing, communication, and memory resources often suffer from stragglers and large variance in per-round latency, which inevitably undermine the overall training efficiency [11].
@@ II-A1 Model Dispatch
The server partitions the pre-trained LLM into an emulator with parameters $\boldsymbol{w}^{E}$ and an adapter with parameters $\boldsymbol{w}^{A}$ [15]. To address the client heterogeneity, the server applies structured pruning to the emulator before transmitting the model to clients, by removing entire attention heads and MLP neurons [12, 6]. Given the heterogeneous client conditions, the server assigns each client $k$ a client-specific pruning rate $\beta_{k}$ and bandwidth $B_{k}$, and then prunes the emulator parameters as
@@
where $\mathcal{P}(\cdot)$ denotes a structured pruning function that removes attention heads and MLP neurons according to the pruning rate $\beta_{k}$. After pruning, the server sends a compressed model package:
@@ II-A3 Parameter Upload
where $\Delta\boldsymbol{w}^{A}_{k}$ represents the parameter delta, namely the difference between the locally updated adapter parameters and the original adapter parameters. Note that no raw data or emulator parameters are transmitted.
@@ II-A4 Update Aggregation
where $|\mathcal{D}_{k}|$ denotes the size (i.e., the number of samples) of the local dataset at client $k$. Then, the server further fine-tunes the emulator on public data, adjusts $\{\beta_{k},B_{k}\}$ for the next round, and redistributes the models.
@@ III Problem Formulation and Optimization
4:  Pruning Update: For each client $k$, fix $\{B_{k}\}$ and update $\beta_{k}$ by
@@
5:  Bandwidth Update: Fix $\{\beta_{k}\}$ and update $\{B_{k}\}$ by solving:
@@
    where $\xi$, $\phi$, and $\psi$ are positive constants, and $\gamma_{\min}$ bounds the gradient norm. This constraint ensures the stability and convergence of the fine-tuning process [8].
@@ IV Experimental Results
Figure 2: The comparison of test loss on Yahoo Answers and GSM8K.
@@ IV-A Experiment Setup
To evaluate the effectiveness of the proposed method, we conduct a series of experiments on various datasets and models.
@@ IV-A2 Fine-tuning Model
We use GPT-2 Medium with 24 decoder layers, wherein layers 0-21 are designated as the emulator and layers 22-23 form the adapter. Client-side fine-tuning is restricted to the adapter, with the emulator remaining frozen. Each client only uploads the updated adapter parameters in each round.
@@ IV-B Results and Discussion
Figs. 2(a) and 2(b) illustrate the effect of federated fine-tuning strategy on the test loss under different total fine-tuning times, with experiments conducted on the Yahoo Answers and GSM8K datasets, respectively (the x-axis denotes total fine-tuning time (s), and the y-axis denotes test loss). First, it is observed that JCPBA consistently achieves the same or even lower test loss in less time compared with all baselines on both datasets. Second, Fig. 2(a) shows that JCPBA attains a test loss of 2.2 at approximately 1,000 s in the Yahoo Answers dataset, whereas UBFP and Split-LoRA-FedFT consume more than 2,000 s to reach a comparable test loss. Meanwhile, JCPBA achieves a lower final test loss than DEFT, FedBiOT, and Split LoRA. On the GSM8K dataset, Fig. 2(b) shows that JCPBA reaches a test loss of 1.1 after 1,500 s, while UBFP needs more than 2,000 s to reach a test loss of around 1.15. In terms of the test loss, JCPBA also outperforms DEFT, FedBiOT, and Split LoRA. This observation corroborates that JCPBA efficiently reduces both computation and communication overhead while preserving and even improving model quality over wireless networks.
@@ IV-B3 Fine-tuning Latency under Heterogeneous Computation Capacities
Fig. 4 compares the fine-tuning latency (total time to reach a fixed target test loss) of the proposed JCPBA with baselines under different levels of computation heterogeneity, where client computation factors $\tilde{f}_{k}$ are sampled from three ranges (i.e., $[1,1.5]$ at low coefficient of variation (CV) of 0.12, $[0.5,2]$ at medium CV of 0.39, $[0.2,2.5]$ at high CV of 0.57). First, we observe that the fine-tuning latency of all methods increases as computation heterogeneity enlarges due to straggler effects, but JCPBA exhibits the smallest latency growth across all heterogeneity levels. Specifically, JCPBA’s latency only rises from 75.2 s at CV of 0.12 to 80.1 s at CV of 0.57, i.e., an increase of about 6%. However, in contrast, UBFP’s latency rises from 96.2 s to 141.3 s (i.e., around 47% growth), FedBiOT from 152.2 s to 250.1 s (i.e., around 64% growth), DEFT from 224.9 s to 279.8 s (i.e., around 24% growth), and Split-LoRA-FedFT from 208.3 s to 574.0 s (i.e., around 176% growth). These results corroborate that JCPBA effectively mitigates straggler effects and sustains efficiency under client heterogeneity.

### v2

# Low-Latency Federated Fine-Tuning for Large Language Models Over Wireless Networks

@@ I Introduction
Although integrating FL and LLMs offers a promising solution to collaboratively fine-tune a shared LLM [15, 19, 2], two major challenges remain. First, the massive scale of LLMs imposes heavy computation and communication demands, which hinders deployment in realistic edge device environments [18]. Second, clients with limited and heterogeneous computing, communication, and memory resources often suffer from stragglers and large variance in per-round latency, which inevitably undermine the overall training efficiency [11].
@@ II-A1 Model Dispatch
The server partitions the pre-trained LLM into an emulator with parameters $\boldsymbol{w}^{E}$ and an adapter with parameters $\boldsymbol{w}^{A}$ [16]. To address the client heterogeneity, the server applies structured pruning to the emulator before transmitting the model to clients, by removing entire attention heads and multi-layer perceptron (MLP) neurons  [12]. Given the heterogeneous client conditions, the server assigns each client $k$ a client-specific pruning rate $\beta_{k}$ and bandwidth $B_{k}$, and then prunes the emulator parameters as
@@
where $\mathcal{P}(\cdot)$ denotes a structured pruning function that removes attention heads and MLP neurons based on the pruning rate $\beta_{k}$. After pruning, the server sends a compressed model package:
@@ II-A3 Parameter Upload
where $\Delta\boldsymbol{w}^{A}_{k}$ represents the parameter delta, namely the difference between the locally updated adapter parameters and the original adapter parameters.
@@ II-A4 Update Aggregation
where $|\mathcal{D}_{k}|$ denotes the size (i.e., the number of samples) of the local dataset at client $k$. After aggregation, the server fine-tunes the emulator once per federated fine-tuning round with public data to maintain the general representation capability while aligning with task-specific patterns. Due to the abundant computational resources at the server, the server-side fine-tuning latency is negligible compared with that of the client-side, thereby not being included in the total fine-tuning latency.
Note that no raw data or emulator parameters are transmitted, which thereby reduces the exposure of sensitive information compared with transmitting full model updates.
@@ III Problem Formulation and Optimization
4:  Pruning Update: Fix $\{B_{k}\}$ and update $\beta_{k}$ by
@@
5:  Bandwidth Update: Fix $\{\beta_{k}\}$ and update $\{B_{k}\}$ by
@@
    where $\xi$, $\phi$, and $\psi$ are positive constants, and $\gamma_{\min}$ bounds the gradient norm.
    The convergence constraint C5 in P1 is used to promote the convergence stability by bounding the gradient norms of federated LLM fine-tuning [15].
@@
It is notable that JCPBA addresses heterogeneous computation and memory capacities by optimizing client-specific pruning rates and bandwidth allocations. The pruning rate is adjusted based on each client’s computation and memory capacity, allowing JCPBA to efficiently adapt to clients with varying resource constraints. This flexibility ensures that JCPBA can maintain more stable performance across heterogeneous environments than the existing related works.
Because the pruning update and the bandwidth update have linear complexity with respect to the number of clients in each iteration, the overall computational complexity of Algorithm 1 is $O(T_{\max}K)$, where $K$ is the number of clients and $T_{\max}$ is the number of iterations, respectively. Since $T_{\max}$ is small in practice, the computational overhead of Algorithm 1 is negligible compared with the client-side LLM fine-tuning cost.
@@ IV-A Experiment Setup
To evaluate the effectiveness of our method, we conduct a series of experiments on various datasets and models.
@@ IV-A2 Fine-tuning Model
We use GPT-2 Medium with 24 decoder layers, wherein layers 0-21 are designated as the emulator and layers 22-23 form the adapter. Client-side fine-tuning is restricted to the adapter, with the emulator remaining frozen. Each client only uploads the updated adapter parameters in each round. We set the learning rate and batch size to $1\times 10^{-4}$ and 4, respectively.
@@ IV-A5 Baselines
Figure 2: The comparison of test loss on Yahoo Answers and GSM8K.
@@ IV-B Results and Discussion
Figs. 2(a) and 2(b) illustrate the effect of federated fine-tuning strategy on the test loss under different total fine-tuning times, with experiments conducted on the Yahoo Answers and GSM8K datasets, respectively (the x-axis denotes total fine-tuning time (s), and the y-axis denotes test loss). First, it is observed that JCPBA consistently achieves the same or even lower test loss in less time compared with all baselines on both datasets. Second, Fig. 2(a) shows that JCPBA attains a test loss of 2.2 at approximately 1,000 s in the Yahoo Answers dataset, whereas UBFP and Split-LoRA-FedFT consume more than 2,000 s to reach a comparable test loss. Meanwhile, JCPBA achieves a lower final test loss than DEFT, FedBiOT, and Split LoRA. On the GSM8K dataset, Fig. 2(b) shows that JCPBA reaches a test loss of 1.1 after 1,500 s, while UBFP needs more than 2,000 s to reach a test loss of around 1.15. In terms of the test loss, JCPBA also outperforms DEFT, FedBiOT, and Split LoRA. This observation corroborates that JCPBA efficiently reduces both computation and communication overhead while preserving and even improving model quality.
@@ IV-B3 Fine-tuning Latency under Heterogeneous Computation Capacities
Fig. 4 compares the fine-tuning latency (total time to reach a fixed target test loss) of the proposed JCPBA with baselines under different levels of computation heterogeneity, where client computation factors $\tilde{f}_{k}$ are sampled from three ranges (i.e., $[1,1.5]$ at low coefficient of variation (CV) of 0.12, $[0.5,2]$ at medium CV of 0.39, $[0.2,2.5]$ at high CV of 0.57). First, we observe that the fine-tuning latency of all methods increases as computation heterogeneity enlarges due to straggler effects, but JCPBA exhibits the smallest latency growth across all heterogeneity levels. Specifically, JCPBA’s latency only rises from 75.2 s at CV of 0.12 to 80.1 s at CV of 0.57, i.e., an increase of about 6%. n contrast, UBFP’s latency rises from 96.2 s to 141.3 s (i.e., around 47% growth), FedBiOT from 152.2 s to 250.1 s (i.e., around 64% growth), DEFT from 224.9 s to 279.8 s (i.e., around 24% growth), and Split-LoRA-FedFT from 208.3 s to 574.0 s (i.e., around 176% growth). These results corroborate that JCPBA effectively mitigates straggler effects and sustains efficiency under client heterogeneity.